Circular No. 6618

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GALACTIC CENTER REGION J. in ’t Zand and J. Heise, Space Research Organization of the Nether-

lands (SRON), Utrecht; A. Bazzano and P. Ubertini, Istituto di Astrofisica Spaziale, CNR, Frascati, Italy; J. M. Muller and M. Smith, SRON and BeppoSAX, Rome, report: “A 3-day exposure with BeppoSAX Wide Field Camera 1 during Mar. 29–Apr. 1 of the galactic-center field reveals the following preliminary results on four weak sources, of which one is not catalogued. The previously unknown x-ray source was detected at 8 mCrab (1.6–26 keV) on Mar. 29. Its position is α = 17^{h}44^{m}42^{s}, δ = −29^{o}16′.9 (equinox 2000.0), with an error-circle radius of 3′; the source has been designated 1SAX J1744–2916. Another transient source was detected on Mar. 30 at α = 17^{h}45^{m}46^{s}, δ = −29^{o}06′.3. The 3′ error-circle radius includes RX J17457–2904. The average intensity was 9 mCrab. The 5σ upper limits for these two sources, for each of the three days when they were not detected, is 4 mCrab. The two transients SAX J1750.8–2900 and GRS 1737–31 that were active one week earlier (IAUC 6597, 6599, 6606) were not detected above a 5σ upper limit of 2 mCrab for the whole observation period.”

GRB 970228 B. Margon and E. W. Deutsch, University of Washington; and D. Q.

Lamb and F. J. Castander, University of Chicago, write: “Observations with the Astrophysical Research Consortium’s 3.5-m telescope at Apache Point on Mar. 3.1 UT of the BeppoSAX WFC1 error circle (IAUC 6572) show an object near the position of the extended object reported by Groot et al. and Metzger et al. (IAUC 6588). Using the astrometric positions of 7 stars (kindly provided by M. Metzger) that lie within ∼ 1′ of the object, we have measured the position of the object as α = 5^{h}01^{m}46^{s}.61 ± 0^{s}.020, δ = +11^{o}46′53′′.4 ± 0′′.15 in the USNO A1.0 astrometric reference frame. The position of the object thus lies within 0′′.3 of the position of the extended object reported by Metzger et al. in the same reference frame. We measure a magnitude b_{J} = 23.3 ± 0.5, relative to the APM b_{J} photometric system, which is ∼ 2 mag brighter than the B = 25.4 reported by Groot et al. for the extended object on Mar. 9.9 UT. We conclude that the object is the optical transient reported by Groot et al. on IAUC 6584. Our observations show that the optical transient, although fading, was still relatively bright 3.0 days after the γ-ray burst, and confirm the close positional coincidence between the optical transient and the extended object reported by Groot et al. and Metzger et al. on IAUC 6588.”

1997 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green